According to Mr. Champagne, the LSU football team has a 65% chance of winning the first game and a 85% chance of winning the second game next season. Using Mr. Champagne's estimates, what is the probability that LSU will LOSE both games? (no ties).
Answer by ikleyn(52781)

.

    It is  P = (1-0.65)*(1-0.85) = 0.35*0.15 = 0.0525.    ANSWER


assuming that the events winning (and respectively loosing) are independent.